OpenStudy (anonymous):

The sum of the fourth and twelfth term of an arithmetic progression is 20. What is the sum of the first 15 terms of the arithmetic progression?

OpenStudy (michele_laino):

we can write these formulas: \[\begin{gathered} {a_4} = {a_1} + 3d \hfill \\ {a_{12}} = {a_1} + 11d \hfill \\ \end{gathered} \] where \(a_1,a_4,a_{12}\) are the first, the fourth and the twelft terms of your progression, and \(d\) is the constant of progression

OpenStudy (michele_laino):

using your condtition we get: \[{a_4} + {a_{12}} = 2{a_1} + 14d = 20\] or: \[{a_1} + 7d = 10\]

OpenStudy (michele_laino):

finally we can write: \[\begin{gathered} {S_{15}} = \frac{{{a_1} + {a_{15}}}}{2} \cdot 15 = \frac{{{a_1} + {a_1} + 14d}}{2} \cdot 15 = \hfill \\ \hfill \\ = \frac{{2{a_1} + 14d}}{2} \cdot 15 = \left( {{a_1} + 7d} \right) \cdot 15 = 10 \cdot 15 = ...? \hfill \\ \end{gathered} \]